# Cortex AISQL: A Production SQL Engine for Unstructured Data

### v1

@@ 3.5.2 Task Specific Aggregation
Algorithmically, is nearly identical to AI_SUMMARIZE_AGG, excepting that each LLM request to LLM.Extract, LLM.Combine, and LLM.Summarize support an additional argument for a user-provided natural language task instruction . For example users may,
1.
Request an aggregation be translated to another language
2.
Identify the three most well-written examples for a product splash page
3.
Write a summary for a technical sales conference instead of a general audience
In practice, the implementation of AI_SUMMARIZE_AGG is simply a curried version of using “Please generate a very concise but informative summary”.
4 Customer Workloads
@@
As we discussed in previous sections, executing semantic operations inside a database engine poses a number of challenges from a system design perspective. At the same time, customers are also adjusting to the AI-centric data management era and are experimenting with different workloads and patterns in order to get the most value out of AI. In this section, we discuss insights from AISQL workloads that were executed in a three month period starting July 2025. We considered actual Snowflake customer workloads across a number of deployments and cloud providers. This workload analysis will hopefully provide some useful insights to both practitioners and researchers.
Our main observations from the workload analysis are as follows:
1.
@@
AI operators consume the bulk of the cost of running AISQL queries. This observation follows from the breakdown of AISQL queries by statement type (see Figure 2) and the proportion of the cost of each query type that comes from LLM GPU credits vs SQL Warehouse credits (see Figure 4). Thus, in the next section, we present optimization techniques that reduce the cost of running AISQL queries with (a) AI-Aware Query Optimization that takes into account the high cost of executing AI operators to create an effective query plan; and (b) Adaptive Model Cascades that uses a lightweight proxy LLM to process most rows while using a larger oracle LLM to handle difficult cases where the proxy model is uncertain.
- 2.
5 AISQL Query Execution Engine
In this section, we discuss the changes we made in Snowflake’s core database engine in order to support SQL queries that use one or more AI operators; for brevity we refer to these queries as AISQL queries. In Section 5.1, we discuss query plan optimizations specific to AISQL queries. In Section 5.2, we discuss how model cascades techniques are used to boost the performance of AI_FILTER. Section 5.3 discusses how AI_JOIN is optimized for classification tasks, and Section 5.4 presents the design and implementation of AI_AGG.
5.1 Optimizing AI Operators
@@
One of the key challenges in optimizing AISQL queries is that AI operators are typically treated as a black box by the optimizer. For example, without historical data from previous invocations on the same data, it is not possible to estimate the selectivity and hardware cost of an AI_FILTER. At the same time, the per-row monetary and runtime cost of AI operators is typically orders of magnitude higher than other typical SQL operators and functions. Consequently, suboptimal placement of an AI operator in the execution plan can have devastating effects both in terms of cost and execution time. This problem is not new in database systems. Optimizing user defined functions (UDFs) in database engines shares many similarities with the problem of optimizing AI operators and many of the techniques that were published 20 years ago are relevant.
@@
Let’s consider the following scenario in which an application like arxiv.org that manages research papers and allows users to search for papers using semantic operations. For simplicity, we assume that the content of each research paper is stored in two relational tables papers and paper_images. The former stores the basic information for each paper like "title" and "authors", whereas the latter stores for each paper information about extracted images and graphs. Both tables have a column of type FILE that links each row with a file stored in blob storage like AWS S3. A simplified schema and some sample data are shown in Figure 6.
@@
Now let’s consider a user who is interested in identifying the research papers that: a) were published in a specific time period, b) discuss "energy efficiency in database systems", and c) include a figure with results using the well known TPC-H benchmark. Finally, the user is interested in compiling a summary of the corresponding abstracts. This query requires inspecting both the textual and the image contents of research papers. Today, the user can write the following AISQL query in Snowflake:
@@
The AISQL query in this example uses two AI functions (AI_SUMMARIZE_AGG and AI_FILTER), has three predicates with different cost and performance characteristics. The first is a predicate on "date" column which is relatively "cheap" to evaluate and for which most database engines can do a reasonably good job at estimating its selectivity. The second predicate is an AI_FILTER on a VARCHAR column. It is more expensive as it invokes an LLM and its selectivity is unknown at compile time. The execution cost can be estimated based on the average number of tokens of the column values. The third predicate is also an AI_FILTER but it is applied on images and is invoking a multimodal LLM. Multimodal LLMs come at a higher cost in terms of hardware resources (add reference). Moreover, the volume of tokens processed for images during inference calls often significantly surpasses that for textual data.
@@
Ignoring the cost of each predicate, the optimizer may generate the Plan A that is shown in Figure 7 in which all predicates are "pushed" below the join, with the underlying assumption that join is an expensive operator and hence the number of rows sent to this operator should be minimized. Also, the most selective predicate (AI_FILTER in this example) is applied first on the rows scanned from the "papers" table. Plan A results in 110,000 LLM calls. The numbers in the parentheses indicate the output cardinality of each operator.
@@
Figure 7 shows a better execution plan (Plan B) that takes into account the cost of the AI predicates and tries to minimize the number of LLM calls. By "pulling" the image AI_FILTER predicate above the join and by changing the order in which predicates are applied on the rows from the "papers" table, the compiler generates an execution plan that results in 330 LLM calls, which results in a 300X improvement in both cost and execution time!
@@
This example highlights two of the optimizations that the Snowflake execution engine is applying in order to improve the performance of AISQL queries (see Section 6 for an experimental evaluation), namely: a) predicate reordering, b) optimizing query plans for AI inference cost. Predicate reordering has both a compiler and an execution component. During compilation, the predicate evaluation order is determined based on the relative cost of each predicate, i.e. the most expensive predicates are evaluated last. During execution, runtime statistics about the cost and selectivity of predicates are collected. These statistics are used to determine at runtime if a different predicate evaluation order is better. For instance, if an AISQL query evaluates two AI_FILTER predicates on text columns from the same input, we can determine at runtime which one is more selective and change the evaluation order so that the more selective AI_FILTER predicate is applied first.
@@
As we discuss in Section 4, AISQL queries typically spend a significant fraction of execution time in AI operators. Consequently, optimizing AISQL queries requires us to rethink the optimization criteria used in the compiler. At Snowflake, we initially considered how AI_FILTER predicates are placed in the execution plan with respect to joins and aggregations as this problem showed up in numerous customer workloads. Even though the selectivity of an AI_FILTER is unknown at compile time, we identified that simply optimizing for the total number AI inference calls produces good results in many cases. However, as AISQL queries become more complex and the compiler estimations less reliable, these kind of optimizations become less effective in practice. To that end, we’re considering dynamic optimization techniques and caching runtime statistics that can help the compiler make more informed decisions in subsequent AISQL queries. In addition to that, we are considering the optimal placement of every AI function and operator, not just AI_FILTER.
5.2 Adaptive Model Cascades
@@
Implementing model cascades in a production database system presents unique challenges. First, the system must learn an effective routing strategy online from streaming data without prior knowledge of the query’s data distribution. Second, the algorithm must operate in a highly concurrent environment, where multiple workers process different portions of the data in parallel. Third, it must honor user-specified cost budgets while maintaining high result quality. Finally, the routing decisions must adapt dynamically as data characteristics change across batches.
Snowflake’s AISQL engine addresses these challenges through an adaptive threshold learning algorithm that reframes cascade routing as a statistical estimation problem. The algorithm uses a sample of oracle evaluations to learn confidence thresholds that partition rows into three regions: accept region where the proxy is confident and predicts positive, reject region where proxy confidence is low and predicts negative, and uncertainty region which might requires invoking oracle and making conservative decisions.
@@
These thresholds are computed using concentration inequalities that provide probabilistic guarantees with limited samples [kang2020approximate]. The algorithm accounts for sampling bias through importance weighting, ensuring threshold estimates are unbiased despite non-uniform sampling.
@@
5.3 Query Rewriting for Semantic Joins
@@
Semantic joins using AI_FILTER enable powerful natural language-based table joining capabilities. For example, let’s consider a scenario where a user has two Snowflake tables: (1) “Reviews” contains product reviews, and (2) “Categories” includes names of product categories. An simplified schema of these two tables along with some sample data are shown in Figure 8.
@@
Now let’s consider an example where a user wants to associate every product review with one or more categories. The user can write the following AISQL queries that joins Reviews with Categories using an AI_FILTER predicate in the ON clause.
@@
Due to the semantic nature of the join condition, traditional efficient join algorithms such as hash-join or sort-merge join that rely on equality predicates are not applicable. The AISQL execution engine must therefore treat the semantic join as a cross join with quadratic complexity, generating all possible row pairs from Reviews and Categories and evaluating AI_FILTER for each pair. Therefore, many queries like this can result in thousands of LLM calls and high inference costs, even for small data sets. More specifically, the above approach requires LLM calls, where and denote the cardinalities of the respective tables.
@@
After analyzing a wide range of customer queries, we discovered that, in many cases, semantic joins are equivalent to a multi-label classification problem. For instance, in the previous example, rather than applying AI_FILTER for every review-category pair, we can leverage the Snowflake AI_CLASSIFY function that can classify a text or image input into one or more user specified labels (see reference). By re-framing the semantic join into a classification problem we instead execute AI_CLASSIFY for every row of table Reviews using the values of column Categories.label as labels, thereby reducing the number of AI inference calls from 24 (4 reviews 6 labels) to just 4 (one for each review), a improvement.
@@
Automatically detecting such cases is not a trivial task. During the compilation of AISQL queries, an AI-based oracle was introduced that examines every semantic join in order to determine if it can be re-written as a multi-label classification problem. The oracle analyzes the user specified natural language prompt, schema metadata (e.g., column and table names), statistics (e.g., number of distinct values), as well as a sample from each input source. Using all that information, the oracle determines during compilation if a semantic join should be converted into a classification task and which input data set contains the labels for the classification. Subsequently, the transformation is performed as a regular rewrite operation by the compiler. The rewritten query is equivalent to the original user query but adds additional operations in order to: a) reduce the number of input labels to each AI_CLASSIFY call, and b) handle cases where the input of each AI_CLASSIFY call fits in the context window of the user specified LLM model.
5.4 AI Aggregation
@@
Earlier, we described an incremental fold strategy for combining intermediate states within a Reduce operation for both and AI_SUMMARIZE_AGG. However, invoking three separate LLM APIs for LLM.Extract, LLM.Combine, and LLM.Summarize introduces significant overhead for small datasets that fit within the model context window. Therefore a simple "short-circuit" alternative provides opportunities for throughput improvements by identifying these scenarios and skipping unnecessary incremental fold steps. This simple algorithm change measured an 86.1% latency reduction in AI_SUMMARIZE_AGG on queries using small data sets.
6 Experimental Evaluation
@@
In this section, we discuss the experimental evaluation for some of the techniques discussed in the previous sections. Due to space constraints, we present a subset of evaluation results, focusing on some of the key techniques that were discussed in the previous sections. All the experiments were conducted using a production-released version of Snowflake.
6.1 AI-aware Query Optimization
@@
In this section, we evaluate the performance impact of two compiler optimizations: a) predicate reordering and b) LLM cost-optimized predicate placement with respect to joins. Figure 9 shows the performance impact of predicate reordering for an AISQL query that is shown below. The SQL statement has two predicates in the WHERE clause, an IN predicate and an AI_FILTER predicate on a VARCHAR column. We used a dataset with 1000 articles from New York Times.
@@
In this experiment, we vary the selectivity of the IN predicate between and , with indicating that all the input rows satisfy the IN predicate. We then measure the speedup we get by carefully reordering the predicate evaluation so that the AI_FILTER predicate is evaluated last. As can be seen from the normalized results in Figure 9, applying the AI_FILTER last results in up to speedup in our experiment. In practice, the actual speedup will depend on a number of factors such as the relative costs of the different predicates and their selectivities.
@@
In this experiment, we vary the ratio between the number of input rows from the left input (), which is where the AI_FILTER is applied, and the total number of output rows from the join. We varied the ratio between and , where means that the join produces twice as much rows as the cardinality of the left input. We compared the optimization that considers the total LLM cost when deciding how to place the AI_FILTER against two other approaches: a) Always Pull-up and, b) Always Push-down. Always Pull-up always "pulls" AI predicates on top of joins, whereas Always Push-down is the default behavior in Snowflake’s optimizer that always "pushes" predicates below joins.
As can be seen from the results in Figure 10, neither Always Pull-up nor Always Push-down are optimal across the different output/input ratios. When the ratio is (i.e. a selective join), Always Pull-up performs better. In contrast, Always Push-down is better when the output/input ratio is (i.e. explosive joins). Considering the total LLM cost when deciding the placement of AI_FILTER predicates (termedAI-aware Optimization) achieves the best performance across the entire range of measured output/input ratios. In practice, the effectiveness of this technique relies on the compiler’s ability to accurately estimate join selectivity, which is not a trivial problem especially in complicated queries with large number of joins.
6.2 Adaptive Model Cascades

### v2

@@ 3.5.1. Task Specific Aggregation
Algorithmically, is nearly identical to AI_SUMMARIZE_AGG, excepting that each LLM request to LLM.Extract, LLM.Combine, and LLM.Summarize support an additional argument for a user-provided natural language task instruction .
The user can provide a more specific task instruction for the aggregation as follows:
3.6. Supporting Multimodal Input
To support multimodal input such as images, audio, or documents, AISQL added a new data type, termed FILE. A FILE value stores a URI as well as various metadata (e.g. size, mime type, creation date, etc) of a file that lives in cloud storage and is managed by the user.
Several utility functions are provided that allow users to manage and process FILEs. For instance, processing only the image files that are stored in a Snowflake table can be performed as follows using the boolean FL_IS_IMAGE function that checks if a file is of a known image file type:
4. Customer Workloads
@@
As discussed earlier, executing semantic operations inside a database engine introduces several system-design challenges. At the same time, customers are adapting to the AI-centric data management era and explore new workloads and usage patterns to get the most value out of AI. This section analyzes AISQL workloads executed over a three-month period (July–September 2025). We considered actual Snowflake customer workloads across multiple deployments and cloud providers. We report observations intended to inform practitioners and motivate further research. Our main observations from the workload analysis are as follows:
@@
AI operators dominate AISQL query cost. This observation follows from the analysis of AISQL query composition by statement type (see Figure 2) and the cost distribution between LLM GPU credits and SQL Warehouse credits for each statement type (see Figure 4). Thus, in the next section, we present optimization techniques that reduce the cost of running AISQL queries with (a) AI-Aware Query Optimization that takes into account the high cost of executing AI operators to create an effective query plan (see Section 5.1); and (b) Adaptive Model Cascades that uses a lightweight proxy LLM to process most rows while using a larger oracle LLM to handle difficult cases where the proxy model is uncertain (see Section 5.2).
Semantic joins are prevalent in customer workloads.
Nearly of AISQL queries involve multiple tables (see Figure 5), and these multi-table queries account for over of total execution time (see Figure 3). Thus, we propose a new method for optimizing semantic joins that reduces the cost and latency of the operation (see Section 5.3).
5. AISQL Query Execution Engine
In this section, we describe the modifications made to Snowflake’s core database engine to enable AISQL queries, i.e. SQL queries that use one or more AI operators. Section 5.1 presents query plan optimizations tailored to AISQL workloads. Section 5.2 discusses the use of model cascade techniques to improve the efficiency of the AI_FILTER operator. Section 5.3 details an optimization that rewrites semantic joins into a multi-label classification problem using AI_CLASSIFY, and Section 5.4 covers enhancements applied to aggregation operators such as AI_AGG and AI_SUMMARIZE_AGG.
5.1. Optimizing AI Operators
@@
One of the key challenges in optimizing AISQL queries is that AI operators are typically treated as black boxes by the optimizer. For example, without historical data from previous invocations on the same data, it is not possible to estimate the selectivity and hardware cost of an AI_FILTER predicate. At the same time, the per-row monetary and runtime costs of AI operators are typically orders of magnitude higher than those of other typical SQL operators and functions. Consequently, the suboptimal placement of an AI operator in the execution plan can have devastating effects on both cost and execution time. This problem is not new in database systems. Optimizing user-defined functions (UDFs) in database engines shares many similarities with the problem of optimizing AI operators, and many of the techniques published 20 years ago remain relevant (hellerstein6caching; hellestein93predicate; chaudhuri99optimization).
@@
To better illustrate the challenges related to optimizing AISQL queries, let us consider the following scenario: a web application like arxiv.org manages research papers and allows users to search for papers using semantic operations. For simplicity, we assume that the content of each research paper is stored in two relational tables papers and paper_images. The former stores the basic information for each paper, such as ”title” and ”authors”, whereas the latter stores information about extracted images and graphs for each paper. Both tables have a column of type FILE that links each row with a file stored in cloud storage (e.g., AWS S3). A simplified schema and some sample data are shown in Figure 6.
@@
Now let us consider a user who is interested in identifying the research papers that: a) were published in a specific time period, b) discuss ”energy efficiency in database systems”, and c) include a figure with results using the well-known TPC-H benchmark. Finally, the user is interested in compiling a summary of the corresponding abstracts. This query requires inspecting both the textual and image contents of research papers. Today, the user can write the following AISQL query in Snowflake as follows:
@@
The AISQL query in this example has three predicates with different cost and performance characteristics. The first is a predicate on a ”date” column, which is relatively ”cheap” to evaluate and for which most database engines can do a reasonably good job at estimating its selectivity. The second predicate is an AI_FILTER on a VARCHAR column, which is more expensive as it invokes an LLM for each row and its selectivity is unknown at compile time. The execution cost can be estimated based on the average number of tokens in the column values. The third predicate is also an AI_FILTER, but it is applied to images and invokes a multimodal model. Multimodal models are typically larger and more expensive than models operating only on text input.
@@
Suppose the cost of each predicate is ignored. In that case, the optimizer may generate Plan A that is shown in Figure 7, in which all predicates are ”pushed” below the join, with the underlying assumption that join is an expensive operator and hence the number of rows sent to this operator should be minimized. In addition, the most selective predicate (AI_FILTER in this example) is applied first on the rows scanned from the ”papers” table. Plan A results in 110,000 LLM calls; the numbers in the parentheses indicate the output cardinality of each operator.
@@
Figure 7 shows a better execution plan (Plan B) that takes into account the cost of AI predicates and tries to minimize the number of LLM calls. By ”pulling” the image AI_FILTER predicate above the join and by changing the order in which predicates are applied on the rows from the ”papers” table, the compiler generates an execution plan that results in 330 LLM calls, which results in a 300 improvement in both cost and execution time!
@@
This example highlights two of the optimizations that the Snowflake execution engine is applying to improve the performance of AISQL queries (see Section 6 for an experimental evaluation), namely: a) predicate reordering, b) optimizing query plans for AI inference cost. Predicate reordering has both a compiler and an execution component. During compilation, the predicate evaluation order is determined based on the relative cost of each predicate, i.e., the most expensive predicates are evaluated last. During execution, runtime statistics are collected on the cost and selectivity of predicates. These statistics are used to determine at runtime whether a different predicate evaluation order is more effective. For instance, if an AISQL query evaluates two AI_FILTER predicates on text columns from the same input, we can determine at runtime which one is more selective and change the evaluation order so that the more selective AI_FILTER predicate is applied first.
@@
As we discussed in Section 4, AISQL queries typically spend a significant fraction of execution time in AI operators. Consequently, optimizing AISQL queries requires us to rethink the optimization criteria used in the compiler. We initially considered how AI_FILTER predicates are placed in the execution plan with respect to joins and aggregations, as this problem showed up in numerous customer workloads. Although the selectivity of an AI_FILTER is unknown at compile time, we identified that simply optimizing the total number of AI inference calls produces good results in many cases. However, as AISQL queries become more complex and compiler estimations less reliable, these kinds of optimizations become less effective in practice. To that end, we are considering dynamic optimization techniques and caching runtime statistics that can help the compiler make more informed decisions in subsequent AISQL queries. Additionally, we are exploring methods to generalize the compiler’s optimization of AI functions and operator placement in complex execution plans.
5.2. Adaptive Model Cascades
@@
Implementing model cascades in a production database system presents unique challenges. First, the system must learn an effective routing strategy online from streaming data without prior knowledge of the query’s data distribution. Second, the algorithm must operate in a highly concurrent environment, where multiple workers process different portions of the data in parallel. Third, it must honor user-specified cost budgets while maintaining high-quality results. Finally, the routing decisions must adapt dynamically as data characteristics change across batches.
Snowflake’s AISQL engine addresses these challenges through an adaptive threshold learning algorithm that reframes cascade routing as a statistical estimation problem. The algorithm uses a sample of oracle evaluations to learn confidence thresholds that partition rows into three regions: an accept region where the proxy is confident and predicts positive, a reject region where proxy confidence is low and predicts negative, and an uncertainty region that might require invoking the oracle and making conservative decisions.
@@
These thresholds are computed using concentration inequalities that provide probabilistic guarantees with limited samples (kang2020approximate). The algorithm accounts for sampling bias through importance weighting, ensuring threshold estimates are unbiased despite non-uniform sampling.
@@
5.3. Query Rewriting for Semantic Joins
@@
Semantic joins using AI_FILTER enable powerful natural language-based table joining capabilities. For example, let us consider a scenario where a user has two Snowflake tables: (1) “Reviews” contains product reviews, and (2) “Categories” includes names of product categories. A simplified schema of these two tables, along with some sample data, is shown in Figure 8.
@@
Consider an example where a user wants to associate every product review with one or more categories. The user can write the following AISQL queries that join Reviews with Categories using an AI_FILTER predicate in the ON clause.
@@
Due to the semantic nature of the join condition, efficient join algorithms such as hash-join or sort-merge join, which rely on equality predicates, are not applicable. The AISQL execution engine must therefore treat the semantic join as a cross join, thereby generating all possible pairs from Reviews and Categories and evaluating AI_FILTER for each pair. Therefore, many queries like this can result in thousands of LLM calls and high inference costs, even for small datasets. More specifically, the above approach requires LLM calls, where and denote the cardinalities of the respective tables.
@@
After analyzing a wide range of customer queries, we discovered that, in many cases, semantic joins are equivalent to a multi-label classification problem. For instance, in the previous example, rather than applying AI_FILTER for every review-category pair, we can leverage the Snowflake AI_CLASSIFY function, which can classify a text or image input into one or more user-specified labels. By re-framing the semantic join into a classification problem we can instead execute AI_CLASSIFY for every row of table Reviews using the values of column Categories.label as labels, thereby reducing the number of AI inference calls from 24 (4 reviews 6 labels) to just 4 (one for each review), a 6 improvement in the example above.
@@
Automatically detecting such cases is not a trivial task. During the compilation of AISQL queries, we introduced an AI-based oracle that examines every semantic join to determine if it can be rewritten as a multi-label classification problem. The oracle analyzes the user-specified natural language prompt, schema metadata (e.g., column and table names), statistics (e.g., number of distinct values), as well as sample values from each input source. Using all that information, the oracle determines during compilation if a semantic join should be converted into a classification task and which input dataset contains the labels for the classification. Subsequently, the transformation is performed as a regular rewrite operation by the compiler. The rewritten query is equivalent to the original user query but adds additional operations to: a) reduce the number of input labels to each AI_CLASSIFY call, and b) handle cases where the input of each AI_CLASSIFY call fits in the context window of the user-specified LLM model.
5.4. AI Aggregation
@@
Earlier, we described an incremental fold strategy for combining intermediate states within a Reduce operation for both and AI_SUMMARIZE_AGG. However, invoking three separate LLM APIs for LLM.Extract, LLM.Combine, and LLM.Summarize introduces significant overhead for small datasets that fit within the model context window. Therefore, a simple ”short-circuit” alternative provides opportunities for throughput improvements by identifying these scenarios and skipping unnecessary incremental fold steps. This simple algorithm change measured an 86.1% latency reduction in AI_SUMMARIZE_AGG on queries using small datasets.
6. Experimental Evaluation
@@
In this section, we present the experimental evaluation of selected techniques introduced in the preceding sections. Due to space constraints, we focus on a representative subset of results that exemplify the key findings and highlight the practical effectiveness of the proposed methods. All experiments were performed using a production-release version of Snowflake.
6.1. AI-aware Query Optimization
@@
In this section, we evaluate the performance impact of two compiler optimizations: a) predicate reordering and b) LLM cost-optimized predicate placement with respect to joins. Figure 9 shows the performance impact of predicate reordering for an AISQL query that is shown below. The SQL statement has two predicates in the WHERE clause, an IN predicate and an AI_FILTER predicate on a VARCHAR column. We used a dataset with 1000 articles from the New York Times111https://www.nytimes.com/.
@@
In this experiment, we vary the selectivity of the IN predicate between and , with indicating that all input rows satisfy the IN predicate. We then measure the speedup we get by carefully reordering the predicate evaluation so that the AI_FILTER predicate is evaluated last. As can be seen from the normalized results in Figure 9, applying the AI_FILTER last yields up to a 7 speedup in our experiment. In practice, the actual speedup will depend on several factors, such as the relative costs of the different predicates and their selectivities.
@@
For this experiment, we adjust the ratio of input rows from the left input (NY_ARTICLES_V1), where the AI_FILTER is applied, to the total number of output rows generated by the join. We varied the ratio between and , where means that the join produces twice as many rows as the cardinality of the left input. We compare the optimization that considers the total LLM cost when deciding the placement of AI_FILTER against two other approaches: a) Always Pull-up and, b) Always Push-down. Always Pull-up always ”pulls” AI predicates on top of joins, whereas Always Push-down is the default behavior in Snowflake’s optimizer that always ”pushes” predicates below joins.
As can be seen from the results in Figure 10, neither Always Pull-up nor Always Push-down is optimal across the different output/input ratios. When the ratio is (i.e., a selective join), Always Pull-up performs better. In contrast, Always Push-down is better when the output/input ratio is (i.e., explosive joins). Considering the total cost of LLM when deciding the placement of AI_FILTER predicates (termed AI-aware Optimization) achieves the best performance across the entire range of measured output/input ratios. In practice, the effectiveness of this technique relies on the compiler’s ability to accurately estimate join selectivity, which is not a trivial problem, especially in complicated queries with a large number of joins. Future efforts will explore adaptive optimization techniques, which involve evaluating various execution plans during compilation while collecting runtime statistics, such as the cost and selectivity of AI operators, from sample data to improve decisions regarding AI operator placement.
6.2. Adaptive Model Cascades